# Hybrid Memetic Search for Electric Vehicle Routing with Time Windows, Simultaneous Pickup-Delivery, and Partial Recharges

### v1

@@ Abstract
With growing environmental concerns, electric vehicles for logistics have gained significant attention within the computational intelligence community in recent years. This work addresses an emerging and significant extension of the electric vehicle routing problem (EVRP), namely EVRP with time windows, simultaneous pickup-delivery, and partial recharges (EVRP-TW-SPD), which has wide real-world applications. We propose a hybrid memetic algorithm (HMA) for solving EVRP-TW-SPD. HMA incorporates two novel components: a parallel-sequential station insertion procedure for handling partial recharges that can better avoid local optima compared to purely sequential insertion, and a cross-domain neighborhood search that explores solution spaces of both electric and non-electric problem domains simultaneously. These components can also be easily applied to various EVRP variants. To bridge the gap between existing benchmarks and real-world scenarios, we introduce a new, large-scale EVRP-TW-SPD benchmark set derived from real-world applications, containing instances with many more customers and charging stations than existing benchmark instances. Extensive experiments demonstrate the significant performance advantages of HMA over existing algorithms across a wide range of problem instances. Both the benchmark set and HMA will be open-sourced to facilitate further research in this area.
@@ I Introduction
Over the past few decades, the vehicle routing problem (VRP) has drawn much attention in the computational intelligence community [1, 2, 3], due to its extensive real-world applications. As VRP research moves forward, a new focus has emerged: the electric vehicle routing problem (EVRP) [4]. This area has gained a lot of interest in recent years, as more and more electric vehicles (EVs) are being used in various sectors [5, 6, 7], such as public transit, home deliveries, postal services, distribution tasks, and emergency power supply. While EVs enable zero-emission logistics, operating a fleet of EVs presents additional challenges compared to operating conventional vehicles: i) the lower energy density of batteries compared to the fuel of combustion engine vehicles, necessitating EVs to perform longer detours to recharge batteries; ii) the limited number of public charging stations; and iii) the substantial time required for battery charging. To address these challenges, Erdogan and Miller-Hooks [8] introduced the first EVRP model that incorporated charging stations, as part of their broader study on the green VRP (G-VRP). Built on their work, subsequent studies have progressively incorporated more constraints into EVRP models to reflect real-world scenarios [9, 10].
This work focuses on an emerging and significant extension of EVRP, namely the electric vehicle routing problem with time windows, simultaneous pickup-delivery, and partial recharges (EVRP-TW-SPD), introduced by Akbay et al. [11]. Compared to the basic EVRP model, EVRP-TW-SPD considers more realistic problem characteristics and constraints, encompassing a wider range of real-world scenarios. Specifically, this model includes constraints such as limited driving range that requires en route recharging and strict customer service time windows [12]. Additionally, it incorporates two key characteristics: reverse logistics and partial recharges. In reverse logistics [13, 14], customers have both pickup and delivery demands, requiring EVs to complete both services in one trip. Partial recharges [15, 16] mean that EVs are not necessarily fully charged each time they visit a charging station, saving time to serve customers with tight time windows.
@@ II Related Work
Reverse logistics, which involves the bi-directional flow of goods in pickup and delivery, is crucial in modern transportation due to its importance in satisfying distinct customer demands and lowering energy consumption. According to [23], the VRPs involving reverse logistics have often been referred to as pickup and delivery problems (PDPs). Among PDPs, the VRP with simultaneous pickup-delivery (VRPSPD) is the most studied variant [24, 25], where vehicles deliver goods from a central depot to customers while also picking up items from customers to return to the depot. VRPSPD was first studied in [26] for a book distribution system with 22 customers and two vehicles, demonstrating that combining pickup and delivery operations leads to significant savings in time and distance compared to separate operations. Since then, numerous studies have been conducted on VRPSPD and its variants. In EVRP-TW-SPD, to handle simultaneous pickup-delivery, existing techniques from VRPSPD research [27, 28, 13] can be adapted. Specifically, in this work we integrate the residual capacity and radical surcharge (RCRS) [13] construction heuristic into the HMA algorithm.
@@
Currently, research on EVRP-TW-SPD is still rather limited. Akbay et al. [11] introduced the problem and formulated it as a mixed integer linear programming (MILP) model. They proposed a metaheuristic approach called Adapt-CMSA-STD and compared it with CPLEX and other heuristics on instances with up to 100 customers and 21 charging stations. Their follow-up work [17] presented Adapt-CMSA-SETCOV, which used vehicle-route representations and a set-covering-based MILP model and showed performance improvement over Adapt-CMSA-STD. Feng et al. [18] developed an alternating direction multiplier method for EVRP-TW-SPD with full recharges, testing it on small-scale instances with up to 15 customers and 6 charging stations.
@@ III Notations and Problem Definition
Formally, the EVRP-TW-SPD is defined on a complete directed graph $\bm{G}=(\bm{V}^{\prime},\bm{E})$, where $\bm{V}^{\prime}=\{0\}\cup\bm{V}\cup\bm{F}=\{0,1,2,\ldots,M+P\}$ is the node set. Here, 00 represents the depot, $\bm{V}=\{1,2,\ldots,M\}$ is a set of $M$ customers, and $\bm{F}=\{M+1,\ldots,M+P\}$ is a set of $P$ charging stations allowing multiple visits. The arc set $\bm{E}=\{(i,j)\mid i,j\in\bm{V}^{\prime},i\neq j\}$ is defined between each pair of nodes. Each node $i\in\bm{V}^{\prime}$ has five attributes: delivery demand $u_{i}$, pickup demand $v_{i}$, time window $[e_{i},l_{i}]$, and service time $s_{i}$. For a customer $i$, $u_{i}$ is the amount of goods delivered from the depot, and $v_{i}$ is the amount of goods picked up to be returned to the depot. The time window $[e_{i},l_{i}]$ defines the earliest and latest time for customer service, with arrivals before $e_{i}$ resulting in waiting and after $l_{i}$ being infeasible. $s_{i}$ is the unloading/loading time at customer $i$. For the depot 00, $e_{0}$ and $l_{0}$ are the earliest departure and latest return time for EVs, respectively, with $u_{0}=v_{0}=s_{0}=0$. For a charging station $i\in\bm{F}$, $u_{i}=v_{i}=s_{i}=0$, and $e_{i}=e_{0}$, $l_{i}=l_{0}$. Each arc $(i,j)\in\bm{E}$ is associated with a travel distance $d_{ij}$ and a travel time $t_{ij}$.
A homogeneous fleet of EVs, all with identical loading capacity $C$, battery capacity $Q$, and dispatching cost $\mu_{1}$, is initially located at the depot. The EVs depart from the depot, serve customers, and finally return to the depot. When an EV visits a charging station, its battery is recharged at a constant rate of $g>0$, i.e., it takes time $g$ to charge one unit of battery energy. Conversely, the constant $h$ represents the EV’s battery energy consumption rate per unit of travel distance for the EV.
@@
[…]
$\quad b_{n_{i,1}}\geq e_{0}\ \text{and}\ a_{n_{i,L_{i}}}\leq l_{0},\ 1\leq i\leq K,$ (5f)
[…]
@@ IV-A Parallel-Sequential Station Insertion (PSSI)
To transform electricity-infeasible routes into high-quality feasible routes, we propose a parallel-sequential station insertion (PSSI) procedure. Compared to existing sequential insertion procedures [15, 21, 22] that typically lead to suboptimal routes, PSSI can better avoid local optima and generate higher-quality routes with acceptable additional computational cost. As shown in Algorithm 1, PSSI leverages two distinct strategies: parallel station insertion (PSI) and sequential station insertion (SSI). PSI (lines 1-1) employs a genetic algorithm to search for optimal station insertions between all consecutive customer pairs simultaneously, determining both the necessity of insertions and the inserted stations. SSI (lines 1-1), on the other hand, inserts charging stations sequentially along the route. Finally, PSSI selects the better route between PSI and SSI as the final outcome (line 1). By combining PSI and SSI, PSSI enables a more comprehensive exploration of the route space, potentially yielding higher-quality routes compared to purely sequential procedures.
@@ IV-A1 Station Selection
For a given problem instance $\bm{G}=(\bm{V}^{\prime},\bm{E})$, we preprocess the charging stations to rank them for insertion between each pair of nodes. The ranking metric is the extra cost induced by insertion, formally defined as $d_{ik}+d_{kj}-d_{ij}$, where $k$ is a potential charging station to be inserted between nodes $i$ and $j$. To balance efficiency and solution quality, only the top $sr\cdot|\bm{F}|$ ranked charging stations are considered for insertion in both PSI and SSI, where $sr\in(0,1]$ is the selection range parameter.
@@
PSI utilizes a genetic algorithm that consists of two main stages: initialization (lines 1-1) and evolution (lines 1-1). In the initialization stage, all charging stations are first removed from the input electricity-infeasible route $\bm{R}$, resulting in a fixed route $\bar{\bm{R}}=(c_{0},c_{1},c_{2},\ldots,c_{L-1},c_{L})$, where $c_{0}=c_{L}=0$ represent the depot, and $c_{1},c_{2},\ldots,c_{L-1}$ are the $L-1$ customers. We define a binary decision variable $X_{j,j+1}$ to indicate whether a charging station is inserted between $c_{j}$ and $c_{j+1}$ (1 for insertion, 0 for no insertion). Thus, the station insertion information is represented as a binary sequence $\bm{x}=(X_{0,1},X_{1,2},\ldots,X_{L-1,L})$. An initial population of $\alpha L$ binary sequences (i.e., the population size is proportional to $L$, with $\alpha$ representing the proportionality parameter) is first generated, where each $X_{j,j+1}$ is randomly set to 1 or 0 with equal probability (line 1). For each sequence, the binary encoding determines whether to insert a station between two consecutive nodes, and we always attempt to insert the highest-ranked station when insertion is indicated. If this insertion violates constraints, we try the next highest-ranked station until a feasible insertion is found or all stations within the selection range have been attempted. This process results in a set of feasible routes $\bm{R}_{1},\bm{R}_{2},\ldots,\bm{R}_{\alpha L}$.
@@
In the evolution stage, the population evolves over $B$ generations. In each generation, two parent sequences are first randomly selected. A child sequence $\bm{x}_{\text{child}}$ is generated using bitwise XOR crossover (denoted by $\oplus$), inheriting unique insertion positions from each parent. Then, mutation is applied to $\bm{x}_{\text{child}}$: each bit has a low chance of flipping (i.e., 2%), followed by a high chance of flipping from 1 to 0 (i.e., 20%) to encourage fewer station insertions. The corresponding route $\bm{R}_{\text{child}}$ is evaluated. Finally, a sequence in the population is replaced with $\bm{x}_{\text{child}}$ if $\bm{R}_{\text{child}}$ is with higher quality (line 1). The best route $\bm{R}_{\text{PSI}}$ among the final population is selected as the outcome of PSI (line 1).
@@
Unlike PSI, SSI progressively inserts stations along the route (lines 1-1). Specifically, for the first node $n_{\text{right}}$ in the route with a negative battery level, it first finds the last charging station or depot visited before this node, denoted as $n_{\text{left}}$. Then, it evaluates all potential station insertions in the path from $n_{\text{left}}$ to $n_{\text{right}}$. Finally, SSI selects the best feasible insertion to perform, i.e., the one that increases the route distance the least (lines 1-1). However, this greedy procedure can be short-sighted, particularly when multiple stations are inserted between two consecutive nodes. To address this limitation, the SSI procedure includes an additional refinement step (line 1). That is, for consecutive stations in the route, we attempt to replace them with stations based on the preprocessed rankings, starting from the highest-ranked to lower-ranked stations. If an improvement is found, the replacement is made. The final outcome of the SSI procedure is $\bm{R}_{\text{SSI}}$.
@@ IV-A2 Charge Amount Calculation
We propose a new procedure tailored to EVRP-TW-SPD for calculating the partial recharge amount, building upon the work of [16, 20]. Our approach considers both the minimal charge amount required and the available slack time, ensuring that the battery state is sufficient to reach the next charging station or depot while satisfying all time window constraints.
We define a path $\bm{P}_{i}=(f_{i},c_{i,1},c_{i,2},\ldots,c_{i,m_{i}},f_{i+1})$ in route $\bm{R}$, where $f_{i}$ and $f_{i+1}$ are two successive charging stations (or the depot) in the route, separated by $m_{i}$ customers $c_{i,j}\ (j=1,...,m_{i})$. Our approach aims to minimize waiting time by utilizing any available slack time for additional charging. The charge amount at $f_{i}$ consists of two parts: the minimal charge $q_{f_{i},0}$ and the additional charge $q_{f_{i},1}$. Specifically, $q_{f_{i},0}$ is calculated as:

### v2

@@ Abstract
With growing environmental concerns, electric vehicles for logistics have gained significant attention within the computational intelligence community in recent years. This work addresses an emerging and significant extension of the electric vehicle routing problem (EVRP), namely EVRP with time windows, simultaneous pickup-delivery, and partial recharges (EVRP-TW-SPD), which has widespread real-world applications. We propose a hybrid memetic algorithm (HMA) for solving EVRP-TW-SPD. HMA incorporates two novel components: a parallel-sequential station insertion (PSSI) procedure for handling partial recharges that can better avoid local optima compared to purely sequential insertion, and a cross-domain neighborhood search (CDNS) that explores solution spaces of both electric and non-electric problem domains simultaneously. These components can also be easily applied to various EVRP variants. To bridge the gap between existing benchmarks and real-world scenarios, we introduce a new, large-scale EVRP-TW-SPD benchmark set derived from real-world applications, containing instances with many more customers and charging stations than existing benchmark instances. Extensive experiments demonstrate the significant performance advantages of HMA over existing algorithms across a wide range of problem instances. Both the benchmark set and HMA are to be made open-source to facilitate further research in this area.
@@ I Introduction
Over the past few decades, the vehicle routing problem (VRP) has attracted much attention in the computational intelligence community [1, 2, 3], due to its extensive real-world applications. As VRP research advances, a new focus has emerged: the electric vehicle routing problem (EVRP) [4]. This area has garnered significant interest in recent years, as electric vehicles (EVs) become increasingly prevalent in various sectors [5, 6, 7], such as public transit, home deliveries, postal services, distribution tasks, and emergency power supply. While EVs enable zero-emission logistics, operating a fleet of EVs presents additional challenges compared to operating conventional vehicles: i) the lower energy density of batteries compared to the fuel of combustion engine vehicles, necessitating EVs to perform detours to recharge batteries; ii) the limited number of public charging stations; and iii) the substantial time required for battery charging. To address these challenges, Erdogan and Miller-Hooks [8] introduced the first EVRP model that incorporated charging stations, as part of their broader study on the green VRP (G-VRP). Built on their work, subsequent studies have progressively incorporated more constraints into EVRP models to reflect real-world scenarios [9, 10, 11, 12].
This work focuses on an emerging and significant extension of EVRP, namely the electric vehicle routing problem with time windows, simultaneous pickup-delivery, and partial recharges (EVRP-TW-SPD), introduced by Akbay et al. [13]. Compared to the basic EVRP model, EVRP-TW-SPD considers more realistic problem characteristics and constraints, encompassing a wider range of real-world scenarios. Specifically, this model includes constraints such as limited driving range that requires recharging along the route and strict customer service time windows [14]. Additionally, it also integrates reverse logistics and partial recharges. In reverse logistics [15, 16], customers have both pickup and delivery demands, requiring EVs to complete both services in one trip. Partial recharges [17, 18] mean that EVs are not necessarily fully charged each time they visit a charging station, saving time to serve customers with tight time windows.
@@ II Related Work
Reverse logistics, which involves the bi-directional flow of goods in pickup and delivery, is crucial in modern transportation due to its importance in satisfying distinct customer demands and lowering energy consumption. According to [25], VRPs involving reverse logistics have often been referred to as pickup and delivery problems (PDPs). Among PDPs, the VRP with simultaneous pickup-delivery (VRPSPD) is the most studied variant [26, 27], where vehicles deliver goods from a central depot to customers while also picking up items from customers to return to the depot. VRPSPD was first studied in [28] for a book distribution system with 22 customers and two vehicles, demonstrating that combining pickup and delivery operations leads to significant savings in time and distance compared to separate operations. Since then, numerous studies have been conducted on VRPSPD and its variants. In EVRP-TW-SPD, to handle simultaneous pickup-delivery, existing techniques from VRPSPD research [29, 30, 15] can be adapted. Specifically, in this work we integrate the residual capacity and radial surcharge (RCRS) [15] construction heuristic into the HMA algorithm.
@@
Currently, research on EVRP-TW-SPD still remains scarce. Akbay et al. [13] introduced the problem and formulated it as a mixed integer linear programming (MILP) model. They proposed a metaheuristic approach called Adapt-CMSA-STD and compared it with CPLEX and other heuristics on instances with up to 100 customers and 21 charging stations. Their follow-up work [19] presented Adapt-CMSA-SETCOV, which used vehicle-route representations and a set-covering-based MILP model and showed performance improvement over Adapt-CMSA-STD. Feng et al. [20] developed an alternating direction multiplier method for EVRP-TW-SPD with full recharges, testing it on small-scale instances with up to 15 customers and 6 charging stations.
@@ III Notations and Problem Definition
Formally, the EVRP-TW-SPD is defined on a complete directed graph $\bm{G}=(\bm{V}^{\prime},\bm{E})$, where $\bm{V}^{\prime}=\{0\}\cup\bm{V}\cup\bm{F}=\{0,1,2,\ldots,M+P\}$ is the node set. Here, 00 represents the depot, $\bm{V}=\{1,2,\ldots,M\}$ is a set of $M$ customers, and $\bm{F}=\{M+1,\ldots,M+P\}$ is a set of $P$ charging stations allowing multiple visits. The arc set $\bm{E}=\{(i,j)\mid i,j\in\bm{V}^{\prime},i\neq j\}$ is defined between each pair of nodes. Each node $i\in\bm{V}^{\prime}$ has five attributes: delivery demand $u_{i}$, pickup demand $v_{i}$, time window $[e_{i},l_{i}]$, and service time $s_{i}$. For a customer $i$, $u_{i}$ is the quantity of goods delivered from the depot, and $v_{i}$ is the amount of goods picked up to be returned to the depot. The time window $[e_{i},l_{i}]$ defines the earliest and latest time for customer service, with arrivals before $e_{i}$ resulting in waiting and after $l_{i}$ being infeasible. $s_{i}$ is the unloading/loading time at customer $i$. For the depot 00, $e_{0}$ and $l_{0}$ are the earliest departure and latest return time for EVs, respectively, with $u_{0}=v_{0}=s_{0}=0$. For a charging station $i\in\bm{F}$, $u_{i}=v_{i}=s_{i}=0$, and $e_{i}=e_{0}$, $l_{i}=l_{0}$. Each arc $(i,j)\in\bm{E}$ is associated with a travel distance $d_{ij}$ and a travel time $t_{ij}$.
A homogeneous fleet of EVs, each with identical loading capacity $C$, battery capacity $Q$, and dispatching cost $\mu_{1}$, is initially located at the depot. The EVs depart from the depot, serve customers, and finally return to the depot. When an EV visits a charging station, its battery is recharged at a constant rate of $g>0$, i.e., it takes time $g$ to charge one unit of battery energy. Conversely, the constant $h$ represents the EV’s battery energy consumption rate per unit of travel distance for the EV.
@@
s.t. $\quad n_{i,1}=n_{i,L_{i}}=0,\ \forall i\in\{1,\dots,K\},$ (5b)
$\quad\sum_{i=1}^{K}\sum_{j=2}^{L_{i}-1}\mathbb{I}[n_{i,j}=z]=1,\ \forall z\in\{1,\dots,M\},$ (5c)
$\quad 0\leq load_{n_{i,j}}\leq C,\ \forall i\in\{1,\dots,K\},\ \forall j\in\{1,\dots,L_{i}\},$ (5d)
$\quad a_{n_{i,j}}\leq l_{n_{i,j}},\ \forall i\in\{1,\dots,K\},\ \forall j\in\{2,\dots,L_{i}\},$ (5e)
$\quad b_{n_{i,1}}\geq e_{0}\ \text{and}\ a_{n_{i,L_{i}}}\leq l_{0},\ \forall i\in\{1,\dots,K\},$ (5f)
$\quad 0\leq y_{n_{i,j}}\leq Y_{n_{i,j}}\leq Q,\ \forall i\in\{1,\dots,K\},\ \forall j\in\{1,\dots,L_{i}\}.$ (5g)
@@ IV-A Parallel-Sequential Station Insertion (PSSI)
To transform electricity-infeasible routes (i.e., routes that do not meet constraint (5g)) into high-quality feasible routes, we propose a parallel-sequential station insertion (PSSI) procedure. Compared to existing sequential insertion procedures [17, 23, 24] that typically lead to suboptimal routes, PSSI can better avoid local optima and generate higher-quality routes with acceptable additional computational cost. As shown in Algorithm 1, PSSI employs two distinct strategies: parallel station insertion (PSI) and sequential station insertion (SSI). PSI (lines 11) employs a genetic algorithm to search for the best station insertions between all consecutive customer pairs simultaneously, determining both the necessity of insertions and the inserted stations. SSI (lines 11), on the other hand, inserts charging stations sequentially along the route. Finally, PSSI selects the better route between PSI and SSI as the final outcome (line 1). By combining PSI and SSI, PSSI enables a more comprehensive exploration of the route space, potentially yielding higher-quality routes compared to purely sequential procedures.
@@ IV-A1 Station Selection
For a given problem instance $\bm{G}=(\bm{V}^{\prime},\bm{E})$, we preprocess the charging stations to rank them for insertion between each pair of nodes. The ranking metric is the extra cost induced by insertion, formally defined as $d_{ik}+d_{kj}-d_{ij}$, where $k$ is a potential charging station to be inserted between nodes $i$ and $j$. The ranking is in ascending order of this extra cost. To balance efficiency and solution quality, only the top $sr\cdot|\bm{F}|$ ranked charging stations are considered for insertion in both PSI and SSI, where $sr\in(0,1]$ is the selection range parameter.
@@
PSI utilizes a genetic algorithm that consists of two main stages: initialization and evolution. In the initialization stage (lines 11), all charging stations are first removed from the input electricity-infeasible route $\bm{R}$, resulting in a fixed route $\bar{\bm{R}}=(c_{0},c_{1},c_{2},\ldots,c_{L-1},c_{L})$, where $c_{0}=c_{L}=0$ represent the depot, and $c_{1},c_{2},\ldots,c_{L-1}$ are the $L-1$ customers. We define a binary decision variable $X_{j,j+1}$ to indicate whether a charging station is inserted between $c_{j}$ and $c_{j+1}$ (1 for insertion, 0 for no insertion). Thus, the station insertion information is represented as a binary sequence $\bm{x}=(X_{0,1},X_{1,2},\ldots,X_{L-1,L})$. An initial population of $\alpha L$ binary sequences (i.e., the population size is proportional to $L$, with $\alpha$ representing the proportionality parameter) is first generated, where each $X_{j,j+1}$ is randomly set to 1 or 0 with equal probability (line 1). For each sequence, the binary encoding determines whether to insert a station between two consecutive nodes, and we always attempt to insert the highest-ranked station when insertion is indicated. If this insertion violates constraints, we try the next highest-ranked station until a feasible insertion is found or all stations within the selection range have been attempted. Since the initial binary value is assigned randomly, it may fail to find a feasible insertion. When this happens, the sequence is regenerated iteratively until a feasible one is found. This process results in a set of feasible routes $\bm{R}_{1},\bm{R}_{2},\ldots,\bm{R}_{\alpha L}$.
@@
In the evolution stage (lines 11), the population evolves over $B$ generations. In each generation, two parent sequences are first randomly selected (line 1). A child sequence $\bm{x}_{\text{child}}$ is generated using bitwise XOR crossover (denoted by $\oplus$), inheriting unique insertion positions from each parent (line 1). Then, mutation is applied to $\bm{x}_{\text{child}}$: each bit has a low chance of flipping (i.e., 2%), followed by a high chance of flipping from 1 to 0 (i.e., 20%) to encourage fewer station insertions (line 1). The corresponding route $\bm{R}_{\text{child}}$ is evaluated and directly discarded if it is infeasible (line 1). Finally, a sequence in the population is replaced with $\bm{x}_{\text{child}}$ if $\bm{R}_{\text{child}}$ is with higher quality (line 1). The best route $\bm{R}_{\text{PSI}}$ among the final population is selected as the outcome of PSI (line 1).
@@
Unlike PSI, SSI progressively inserts stations along the route (lines 11). Specifically, for the first node $n_{\text{right}}$ in the route with a negative battery level, it first finds the last charging station or depot visited before this node, denoted as $n_{\text{left}}$. Then, it evaluates all potential station insertions in the path from $n_{\text{left}}$ to $n_{\text{right}}$. Finally, SSI selects the best feasible insertion to perform, i.e., the one that increases the route distance the least (lines 11). However, this greedy procedure can be short-sighted, particularly when multiple stations are inserted between two consecutive nodes. Among these multiple stations, the station inserted earlier may no longer be the best choice. To address this limitation, the SSI procedure includes an additional refinement step (line 1). That is, for consecutive stations in the route, we attempt to replace them with stations based on the preprocessed rankings again, starting from the highest-ranked to lower-ranked stations. If an improvement is found, the replacement is made. The final outcome of the SSI procedure is $\bm{R}_{\text{SSI}}$.
@@ IV-A2 Charge Amount Calculation
We propose a new procedure tailored to EVRP-TW-SPD for calculating the partial recharge amount, building upon the work of [18, 22]. Our approach considers both the minimal required charge amount and the available slack time. This ensures that the battery state is sufficient to reach the next charging station (or the depot) while satisfying time window constraints. Moreover, it aims to minimize waiting time by utilizing any available slack time for additional charging.
We define a path $\bm{P}_{i}=(f_{i},c_{i,1},c_{i,2},\ldots,c_{i,m_{i}},f_{i+1})$ in route $\bm{R}$, where $f_{i}$ and $f_{i+1}$ are two successive charging stations (or the depot) in the route, separated by $m_{i}$ customers $c_{i,j}\ (j=1,...,m_{i})$. The charge amount at $f_{i}$ consists of two parts: the minimal charge $q_{f_{i},0}$ and the additional charge $q_{f_{i},1}$. Specifically, $q_{f_{i},0}$ is calculated as: